Max invested $15,000 in two different funds. One earned 6% interest in the first year and the other earned 8%, for a total of $1100 in interest. Determine how much he invested in each fund.
Answer by Boreal(15235)

x at 6%
15000-x at 8%
.06x+.08(15000-x)=1100
.06x+1200-.08x=1100
-.02x=-100
x=5000@ 6% ($300)
15000-x=10000@8%=($800)
total is $1100.